OpenStudy (anonymous):

write the standard form of the equation of the line through the pair points (-3,-1) and (-10, 4) Write the equation using only integer coefficients

OpenStudy (anonymous):

y=(-5x)/7 - (22/7)

OpenStudy (anonymous):

The equation is ?x+?y+?

OpenStudy (anonymous):

5x + 7y =-22 (ax+by =c - no fractional coefficients, a can't be negative also )